OpenStudy (anonymous):

solve the radical expression square root of 5x-9=16

OpenStudy (anonymous):

x=53

OpenStudy (radar):

\[\sqrt{5x-9}=16\] Square both sides 5x-19=256 Add 19 to both sides 5x=275 Divide both sides by 5 x=55

OpenStudy (radar):

Any questions or comments?

OpenStudy (anonymous):

yes and u r wrong its not 19 its 9

OpenStudy (radar):

Ah so, did goof then it would of been: 5x=265 x=53, just as you posted with out showing steps or procedure, Just providing an answer doesn't help learning.

OpenStudy (anonymous):

@ radar if u want to learn then come to twiddla

OpenStudy (radar):

I do learn a lot from the posts, but again, I am not interested in the "answer" but the steps/procedures used to obtain the answer. Curious minds need to know.

OpenStudy (radar):

I do learn from the posts, but as I said, only if the problem is worked out showing the procedures required to get the solution. The answer is of no consequence. Some do this quite well, satellite, and amistre to just name a couple. You will note that they are often requested (by name) to solve problems.